### v1

# Facilitating Long Context Understanding via Supervised Chain-of-Thought Reasoning

Abstract
Recent advances in Large Language Models (LLMs) have enabled them to process increasingly longer sequences, ranging from 2K to 2M tokens and even beyond. However, simply extending the input sequence length does not necessarily lead to effective long-context understanding. In this study, we integrate Chain-of-Thought (CoT) reasoning into LLMs in a supervised manner to facilitate effective long-context understanding. To achieve this, we introduce LongFinanceQA, a synthetic dataset in the financial domain designed to improve long-context reasoning. Unlike existing long-context synthetic data, LongFinanceQA includes intermediate CoT reasoning before the final conclusion, which encourages LLMs to perform explicit reasoning, improving accuracy and interpretability in long-context understanding. To generate synthetic CoT reasoning, we propose Property-driven Agentic Inference (PAI), an agentic framework that simulates human-like reasoning steps, including property extraction, retrieval, and summarization. We evaluate PAI’s reasoning capabilities by assessing GPT-4o-mini w/ PAI on the Loong benchmark, outperforming standard GPT-4o-mini by 20.0%. Furthermore, we fine-tune LLaMA-3.1-8B-Instruct on LongFinanceQA, achieving a 24.6% gain on Loong’s financial subset.
Facilitating Long Context Understanding via
Supervised Chain-of-Thought Reasoning
Jingyang Lin1,2 Andy Wong2, Tian Xia2, Shenghua He2, Hui Wei2,3, Mei Han2, Jiebo Luo1 1 University of Rochester, NY, United States 2 PAII Inc., CA, United States 3 University of California, Merced, CA, United States
footnotetext: Corresponding author: ✉ jlin81@ur.rochester.edu.footnotetext: Work was done during the internship at PAII Inc.
1 Introduction
Figure 1: The hype of long-context large language models. The results shown are from the Loong benchmark, where green points refer to open-source LLMs and red points indicate closed-source LLMs. Our GPT-4o-mini w/ PAI stands out as the red open circle.
Long context understanding remains an evolving challenge Kočiský et al. (2018); Wu et al. (2021); Bai et al. (2024); Wang et al. (2024) in natural language processing (NLP). Achieving long-context understanding requires processing long-form textual information, thereby enhancing a model’s ability to generate coherent, accurate, and contextually relevant responses IBM Research (2024). Practical long context understanding has a potential impact on numerous applications, such as private document analysis Mukherjee et al. (2023), large codebase understanding Nam et al. (2024), and multimodal content understanding Chandrasegaran et al. (2024); Lin et al. (2023); Tang et al. (2023); Chen et al. (2024c). Recent advancements in large language models Ouyang et al. (2022); Reid et al. (2024); Dubey et al. (2024) have significantly extended the input sequence length, ranging from 2K to 2M tokens, as shown in Figure 5 in Appendix A.
However, simply increasing the input sequence length does not necessarily improve the ability to comprehend the long content Yang et al. (2024a); Goldman et al. (2024). In particular, Figure 1 presents the performance of various long-context LLMs on the Loong benchmark Wang et al. (2024). The results suggest that, regardless of the maximum sequence length models can process, their performance remains similar, typically between 45% and 55%. This phenomenon exposes a hype that despite considerably enlarging context window size, the state-of-the-art LLMs still fail to perform satisfactorily in practical long-context problem-solving tasks. Therefore, instead of merely increasing the input sequence length, achieving effective long-context understanding remains an open challenge.
In parallel with advanced long-context architectures and techniques Peng et al. (2024); Liu et al. (2024c); Dao and Gu (2024), constructing high-quality long-context training data remains essential yet underexplored. Given the scarcity and high annotation costs of long-context data, generating high-quality synthetic data for long-context modeling is valuable and urgent. Early attempts Raffel et al. (2020); Fu et al. (2024) simply pack all short-length data into long-context chunks without considering document boundaries. Later works Zhang et al. (2024a); He et al. (2024) have introduced long-context QA tasks from both single- and multi-source perspectives. Qwen-Agent Yang et al. (2024a) develops an agentic system to further improve the quality of synthetic answers.
However, existing synthetic long-context data typically pair challenging questions with brief final answers for model training. This way overlooks a key difference between long-context and traditional QA tasks: practical long-context questions often require multi-step reasoning throughout the long content. Without intermediate reasoning, LLMs struggle to learn effective patterns from the paired complex questions and brief answers. We hypothesize that directly guiding models to generate brief answers without intermediate reasoning steps for long-context modeling will lead to suboptimal training. Instead, incorporating intermediate reasoning into synthetic data will help LLMs learn effective patterns and enhance training optimization.
To validate this hypothesis, we introduce LongFinanceQA, a novel long-context synthetic dataset constructed using financial data. Each sample in this dataset consists of a practical long-context question and the corresponding augmented answer along with intermediate chain-of-thought (CoT) reasoning steps. In particular, we first collect 6,911 bilingual financial annual reports (i.e., English and Chinese) published before 2022. We then build a financial metric pool comprising key metrics commonly found in these reports, such as profit, cash flow, and debt. Based on these financial metrics, we generate 46,457 long-context questions that require single- or multi-source evidence. To generate reliable reasoning-augmented answers, we propose Property-driven Agentic Inference (PAI), a comprehensive agentic framework. PAI leverages LLM-based agents to simulate human-like reasoning and operates in three steps: 1) a property extraction agent decomposes complex queries by identifying key properties, where each property consists of a metric and the corresponding subject; 2) a property-based retrieval agent retrieves relevant information from long documents for each identified property, and then generates intermediate findings by leveraging the property-based retrieved content; 3) a summarization agent synthesizes the intermediate findings into a coherent conclusion. The synthetic reasoning results are formed by integrating outputs from the property extraction and retrieval stages, while the conclusion is derived from the summarization stage. Finally, LongFinanceQA comprises 46,457 long-context QA pairs with CoT reasoning over 6,911 financial reports.
Although PAI performs human-like reasoning in long-context scenarios, it relies on human-crafted design and multi-step inference. To simplify this inference process, we intend to transfer PAI’s long-context reasoning ability to a large language model, LLaMA-3.1, via supervised fine-tuning on LongFinanceQA. The enhanced model, LongPAI, leverages CoT reasoning to handle long-context problems in a single step. This fine-tuning procedure is termed as Supervised CoT Reasoning.
Empirically, we first evaluate the outcome of PAI on the Loong benchmark Wang et al. (2024), involving challenging long-context tasks on three different domains. The results show that equipping GPT-4o-mini with PAI achieves a substantial 20% improvement on the Loong as shown in Figure 1. Moreover, the effectiveness of PAI guarantees the quality of the synthetic data in LongFinanceQA. Meanwhile, the enhanced LLaMA-3.1 model, LongPAI, achieves a 24.6% improvement on the Financial subset of Loong. Notably, in several scenarios, LongPAI even surpasses its teacher model PAI. This phenomenon emphasizes the importance of long-context modeling, contradicting recent arguments that the long-context problem can be solved by short language models Qian et al. (2024); Chen et al. (2024b).
Our main contributions are three-fold: 1) we introduce LongFinanceQA, a long-context synthetic dataset for fine-tuning with 46,457 QA pairs featuring high-quality CoT reasoning from 6,911 bilingual financial annual reports; 2) to generate reasoning-augmented answers, we propose an agentic framework, Property-driven Agentic Inference, to mimic human behaviors; and 3) empirical results validate the effectiveness of PAI and supervised CoT reasoning, the quality of LongFinanceQA, and the importance of long-context modeling.
2 Related Work
Long-context Synthetic Data. The scarcity of well-annotated long-context data makes high-quality synthetic data generation a valuable research direction. Early works Raffel et al. (2020); Fu et al. (2024) concatenate short data into long-context fragments without considering the boundaries of the document. Large World Model Liu et al. (2024b) addresses this boundary issue with masked sequence packing to keep attention within documents. Subsequent studies Zhang et al. (2024a); He et al. (2024) construct single- and multi-source long-context QA pairs requiring evidence retrieval across multiple document positions. Qwen2-Agent Yang et al. (2024a) leverages multiple agents to enhance answer quality. Beyond conventional QA pairs, our study augments answers with intermediate reasoning steps, explicitly guiding language models in learning reasoning abilities for practical long-context scenarios.
Long-context Large Language Models. Two main approaches enhance long-context problem-solving: reduction-based and extension-based methods. Reduction-based methods compress input by preserving essential information, allowing LLMs to focus on relevant content. Techniques like Retrieval-Augmented Generation (RAG) Lewis et al. (2020) and task decomposition for book summarizationWu et al. (2021) follow this approach. Extension-based methods expand the context window directly. RoPE Su et al. (2024) introduces a foundational positional encoding widely used in long-context LLMs Peng et al. (2024); Yang et al. (2024a); Dubey et al. (2024). Parallelism techniques Ren et al. (2021); Liu et al. (2024c) scale context capacity, enabling fully fine-tuning for long-context LLMs. This study leverages parallelism techniques to achieve long-context training.
Chain-of-Thought Resoning. The coT technique improves the reasoning abilities of language models by incorporating intermediate reasoning steps. Early works Wei et al. (2022) present that prompting LLMs with step-by-step reasoning significantly improves performance on complex reasoning tasks. Following works Zelikman et al. (2022); Yao et al. (2024) explore structured CoT approaches, such as tree-based and self-consistent reasoning. This work incorporates CoT reasoning steps into fine-tuning data instead of well-crafted prompts. This approach enables the augmented data to explicitly supervise language models in reasoning skills.
3 Methodology
In this section, we first introduce the problem formulation of the QA task enhanced by intermediate CoT reasoning in Section 3.1. Then, we present the procedure of LongFinanceQA dataset construction in Section 3.2. After that, we describe the fine-tuning details in Section 3.3.
3.1 Problem Formulation
Traditional QA tasks require language models to generate outputs $\mathbf{A}$ directly from a given query $\mathbf{Q}$ and the corresponding input content $\mathbf{X}$ by modeling the conditional probability:
$p_{\theta}(\mathbf{A}|\mathbf{X},\mathbf{Q}),$ (1)
where $\theta$ is the parameters of language models.
Compared to traditional QA tasks, practical long-context QA tasks often require intermediate reasoning steps to analyze multiple pieces of evidence across long documents, and then derive the final answer. Thus, we formalize long-context QA tasks as a joint conditional probability of intermediate reasoning results $\mathbf{R}$ and the final answer $\mathbf{A}$:
$p_{\theta}(\mathbf{R},\mathbf{A}|\mathbf{X},\mathbf{Q}).$ (2)
As shown in Eq.(2), this study aims to facilitate long-context understanding by guiding LLMs to first predict intermediate CoT reasoning steps before generating the final answer. To achieve this, we construct a long-context synthetic dataset that explicitly supervises language models in learning intermediate reasoning. The data construction process can be factorized as follows:
$p_{\theta}(\mathbf{R},\mathbf{A}|\mathbf{X},\mathbf{Q})=p_{\theta}(\mathbf{R}|\mathbf{X},\mathbf{Q})\cdot p_{\theta}(\mathbf{A}|\mathbf{X},\mathbf{Q},\mathbf{R}),$ (3)
where $p_{\theta}(\mathbf{R}|\mathbf{X},\mathbf{Q})$ indicates the generation of intermediate reasoning steps $\mathbf{R}$ given a query $\mathbf{Q}$ and input content $\mathbf{X}$, and then $p_{\theta}(\mathbf{A}|\mathbf{X},\mathbf{Q},\mathbf{R})$ is the process of incorporating the query and summarizing reasoning steps to produce the answer $\mathbf{A}$. The data construction follows the principle of Eq.(3).
[…]
3.2 LongFinanceQA Dataset
LongFinanceQA dataset is designed to generate practical long-context QA pairs with reasoning steps to effectively analyze long content. The finance domain was chosen for several reasons. First, annual financial reports are readily accessible and present complex long-context reasoning challenges. Moreover, finance is a data-driven field where accurate insights can drive critical decisions, making advancements in AI particularly valuable for real-world applications. We will introduce the data construction pipeline of LongFinanceQA as follows.
Data Collection. To begin, we collect bilingual financial annual reports (i.e., English and Chinese) dated before 2022 from open-source official websites, specifically the SEC-10-K [^1] https://www.sec.gov/ and cninfo [^2] http://www.cninfo.com.cn/ platforms. Documents from companies included in the Loong benchmark are then excluded. Next, we filter reports based on a token length range of 20K to 80K, determined using the GPT-4o tokenizer. Moreover, we prioritize companies with consistent annual reports over the years. In the end, 6,911 bilingual financial reports are selected.
Diverse Query Generation. We first construct a financial metric pool containing key metrics commonly found in financial reports, such as profit, revenue, and cash flow. Then, we randomly pick several metrics in the metric pool and select a combination of financial reports. Given these financial metrics and the metadata from the selected documents (e.g., company name and year), we generate various long-context questions that require either single- or multi-source evidence. Next, we filter out questions whose corresponding combined documents exceed 256K tokens, as this surpasses the maximum token limit of our model. Finally, we obtain 46,457 practical long-context questions. Please refer to Appendix A for more details.
Augmented Answer Generation with CoT Reasoning. Given long-context questions and their corresponding documents, our goal is to generate answers with CoT reasoning, following the principle of Eq. (3). In particular, we first generate step-by-step reasoning based on the question and documents, then integrate the reasoning steps into a conclusion. Generally, long-context questions are challenging as they require models to retrieve multiple pieces of evidence scattered throughout long content and then integrate them for a global understanding Wang et al. (2024); Edge et al. (2024). Inspired by this phenomenon, we intend to first extract key evidence points that support answering the question, then retrieve relevant information based on these points, and finally aggregate them into the conclusion. We term these supporting evidence points as “Properties”. Following this methodology, we propose the Property-driven Agentic Inference (PAI) framework, illustrated in Figure 2, containing three steps: (1) property extraction, (2) property-based retrieval, and (3) summarization.
[…]
Step 3: Summarization. In this final step, the original query $\mathbf{Q}$ and all intermediate findings from the second step are integrated to generate a comprehensive conclusion $\mathbf{A}$.
Figure 3: Token length distribution of answers with and without CoT reasoning from Multi-Source QA pairs in the proposed LongFinanceQA dataset.
The PAI serves as a low-cost annotator for the LongFinanceQA dataset. Using PAI, we produce augmented answers with CoT reasoning by combining the intermediate reasoning results from the first two steps with the conclusion. Figure 3 illustrates the token length difference between answers with and without CoT reasoning, showing that reasoning-augmented answers are nearly 200 tokens longer on average. [^3] The answers are tokenized by the LLaMA-3.1 tokenizer. Please check out more data statistics in the Appendix A.
3.3 Supervised CoT Reasoning
Although PAI can behave like a human in long-context scenarios, it requires human-crafted design and multi-step inference. To address this limitation, we seek to transfer the long-context reasoning capability of PAI to existing language models, enabling them to analyze long content in a single-step inference. To this end, we fine-tune a language model on LongFinanceQA to explicitly guide it in learning CoT reasoning. This fine-tuning procedure is termed as Supervised CoT Reasoning. Specifically, we use LLaMA-3.1-8B-Instruct Dubey et al. (2024) as the base language model. Following Fu et al. (2024), we first extend the context window of LLaMA-3.1 from 128K to 262K through continued pretraining on 1.6B tokens, which consist of packed documents from Slimpajama Soboleva et al. (2023). After extending context length, we fine-tune the extended LLaMA-3.1 model (parameterized by $\theta$) on LongFinanceQA with reasoning-augmented answers.
Through this process, we obtain an enhanced LongPAI model by maximizing the log-likelihood of the predicted answer $\mathbf{Y}$ along with the CoT reasoning $\mathbf{R}$, conditioned on the given long documents $\mathbf{X}$ and the query $\mathbf{Q}$. The objective is calculated as:
$\mathcal{L(\theta)}=\sum_{i=1}^{N_{r}+N_{y}}\log p_{\theta}(\mathcal{\mathbf{R},\mathbf{Y}|\mathbf{X},\mathbf{Q}}),$ (4)
where $N_{r}$ is the number of reasoning tokens and $N_{y}$ indicates the answering token, including those in the properties, sub-answers, and final answers. We mask non-answer positions during the fine-tuning, allowing for more efficient learning. Although we only consider the LLaMA-3.1 model here, LongFinanceQA can also be used to fine-tune other language models. Please refer to Section 4.1 for more implementation details.
4 Experiments
Table 1: Data quality assessment for the long-context synthetic dataset (LongFinanceQA) by measuring the performance of PAI on the Loong benchmark. AS denotes Average Scores (0-100), and PR represents the Perfect Rate (0-1). Green highlights the remarkable improvements over the base model (GPT-4o-mini).
Model Context Spotlight Locating Comparison Clustering Chain of Reasoning Overall
Length AS PR AS PR AS PR AS PR AS PR
Open-Source Long-Context Large Language Models
LLaMA-3.1-8B-Instruct 128K 62.42 0.52 39.13 0.21 25.96 0.01 44.20 0.22 38.79 0.18
DeepSeek-R1-Qwen-32B 128K 51.68 0.41 49.25 0.34 41.53 0.16 45.00 0.30 45.45 0.27
Qwen2-72B-Instruct 128K 54.17 0.36 42.38 0.20 36.71 0.04 47.76 0.18 43.29 0.15
Qwen2.5-72B-Instruct 128K 65.08 0.55 51.90 0.30 46.07 0.08 64.43 0.40 54.83 0.28
Qwen2.5-14B-Instruct-1M 1000K 67.50 0.58 55.12 0.35 39.05 0.04 57.81 0.31 51.30 0.25
Closed-Source Long-Context Large Language Models
Kimi-Chat 200K 60.98 0.50 34.74 0.13 28.76 0.04 38.52 0.15 37.49 0.16
Claude3.5-Sonnet 200K 58.45 0.49 54.21 0.35 45.77 0.07 43.92 0.25 48.85 0.23
GPT-4o 128K 73.95 0.62 50.50 0.28 44.29 0.09 57.95 0.28 53.47 0.26
Gemini-1.5-pro 1000K 75.02 0.56 49.94 0.27 44.10 0.09 64.97 0.37 55.37 0.27
GPT-4o-mini (Base) 128K 59.46 0.49 51.90 0.27 34.55 0.04 64.28 0.39 49.25 0.24
GPT-4o-mini w/ PAI (ours) 128K 79.74 0.67 67.60 0.46 62.80 0.27 75.46 0.57 69.58 0.44
Table 2: Performance on Financial subset of Loong benchmark. AS represents Avg Scores (0~100) and PR denotes Perfect Rate (0~1). Green highlights improvements over the base model. Full results are shown in Appendix A.
Model Context Spotlight Locating Comparison Clustering Chain of Reasoning Overall
Length AS PR AS PR AS PR AS PR AS PR
$\mathtt{All\ Set}$ (10K-250K)
Qwen2-72B-Instruct 128K 59.80 0.47 61.12 0.43 34.32 0.06 74.68 0.50 53.20 0.32
Qwen2.5-72B-Instruct 128K 71.07 0.63 59.14 0.41 38.23 0.08 81.09 0.60 57.36 0.36
Qwen-2.5-14B-Instruct-1M 1000K 78.83 0.72 65.27 0.50 36.24 0.07 79.10 0.64 59.78 0.41
GPT-4o 128K 88.23 0.84 62.90 0.48 45.51 0.17 69.40 0.43 63.05 0.44
GPT-4o-mini (Base) 128K 70.90 0.59 59.37 0.38 36.33 0.06 79.58 0.58 56.50 0.34
GPT-4o-mini w/ PAI (ours) 128K 91.07 0.83 74.40 0.58 61.55 0.32 89.63 0.78 75.56 0.57
LLaMA-3.1-8B-Instruct (Base) 128K 67.84 0.56 47.12 0.30 24.62 0.02 63.63 0.34 45.88 0.26
LongPAI (ours) 262K 87.02 0.78 64.50 0.48 60.95 0.32 81.89 0.71 70.54 0.52
$\mathtt{Set3}$ (100K-200K)
Qwen2-72B-Instruct 128K 47.00 0.33 48.07 0.27 25.79 0.00 69.37 0.34 42.98 0.20
Qwen2.5-72B-Instruct 128K 60.47 0.48 49.00 0.28 30.61 0.01 76.54 0.46 48.99 0.26
Qwen-2.5-14B-Instruct-1M 1000K 74.33 0.68 54.64 0.35 30.72 0.01 73.71 0.51 53.47 0.33
GPT-4o 128K 87.25 0.83 46.00 0.31 36.68 0.08 64.57 0.40 54.79 0.36
GPT-4o-mini (Base) 128K 63.05 0.53 53.48 0.24 29.80 0.01 72.37 0.46 50.03 0.26
GPT-4o-mini w/ PAI (ours) 128K 94.08 0.83 74.13 0.63 55.78 0.20 87.71 0.77 74.21 0.55
LLaMA-3.1-8B-Instruct (Base) 128K 63.38 0.47 36.04 0.19 20.28 0.00 62.49 0.26 40.45 0.20
LongPAI (ours) 262K 88.00 0.80 50.07 0.28 55.01 0.21 84.03 0.69 65.10 0.43
$\mathtt{Set4}$ (200K-250K)
Qwen2-72B-Instruct 128K 41.85 0.19 39.75 0.15 29.17 0.03 41.67 0.07 37.23 0.11
Qwen2.5-72B-Instruct 128K 57.48 0.44 49.50 0.30 27.33 0.00 55.00 0.13 45.51 0.22
Qwen-2.5-14B-Instruct-1M 1000K 59.44 0.41 37.00 0.20 27.33 0.00 31.67 0.00 39.57 0.16
GPT-4o 128K 69.26 0.56 50.50 0.35 30.70 0.00 50.67 0.07 49.58 0.25
GPT-4o-mini (Base) 128K 48.37 0.26 50.00 0.30 28.70 0.00 48.33 0.07 42.30 0.15
GPT-4o-mini w/ PAI (ours) 128K 82.78 0.70 63.50 0.35 48.00 0.17 76.00 0.53 66.14 0.42
LLaMA-3.1-8B-Instruct (Base) 128K 40.74 0.30 35.85 0.20 19.77 0.00 28.73 0.00 30.88 0.13
LongPAI (ours) 262K 77.89 0.63 62.00 0.45 48.00 0.17 51.33 0.47 60.36 0.41
4.1 Experimental Setup
[…]
Base Models. We adopt GPT-4o-mini Achiam et al. (2023) and LLaMA-3.1-8B-Instruct Dubey et al. (2024) as our base models. Specifically, GPT-4o-mini serves as the agent of the proposed Preference Agentic Inference (PAI), while LLaMA-3.1-8B is used as the base model for LongPAI, which is fine-tuned on the LongFinanceQA dataset.
Implementation Details. To enable training on long sequences (> 250K), we employ several optimization techniques, including flash-attention-2 Dao (2023) and ring-attention Liu et al. (2024c). Furthermore, we adopt a zigzag sharding approach Zhu (2024) within ring attention for more effective load distribution across multiple GPUs. This training setup allows us to fine-tune the large language models fully. Using 8 A100 GPUs, the long-context training is completed in three days for fine-tuning. In addition, in training the long-context LLMs, we adopt the rotary base scaling approach Liu et al. (2024d) and scale up the base value (i.e., rotary base of 1,247,820) to adapt RoPE to a longer context. For optimization, we use a constant learning rate of 2e-6 for the entire training procedure. Following the common practice Fu et al. (2024), we set the batch size to 16M tokens as mentioned in Dubey et al. (2024). During inference, we set the temperature as zero to eliminate the randomness. We also increase the maximum output tokens to 1,024 since the CoT reasoning requires more output tokens.
4.2 Main Results
In this section, we first assess the quality of our long-context synthetic data (LongFinanceQA) by evaluating the performance of the proposed PAI on the Loong benchmark. Next, we compare the enhanced long-context language model (LongPAI) with its base LLaMA-3.1 model and other state-of-the-art LLMs on the Finance subset of Loong.
Data Quality Assessment. Reasoning-augmented answers in LongFinanceQA are automatically generated by the PAI framework. To assess the quality of these synthetic answers, we evaluate the annotator, PAI, on the Long benchmark, using GPT-4o-mini as the agent within the PAI, referred to as GPT-4o-mini w/ PAI. Table 1 shows that the GPT-4o-mini-based PAI framework significantly enhances the base model’s performance, demonstrating the effectiveness of the PAI. In particular, compared to the standard GPT-4o-mini model, the overall average score improves by 20.3%, with substantial gains in key tasks such as spotlight locating (+20.2%), comparison (+15.7%), clustering (+29.2%), and chain of reasoning (+10.2%). Although the basic GPT-4o-mini is not the strongest model, GPT-4o-mini w/ PAI outperforms the state-of-the-art closed-source model, Geneni-1.5-pro Reid et al. (2024), by over 15%. Moreover, the superior performance in the Spotlight Locating task (i.e., single-source QA task) highlights the quality of intermediate reasoning results generated by the PAI, as predictions in this task serve as essential reasoning steps for the other three multi-source QA tasks. Consequently, the strong performance on the Loong demonstrates the capability of PAI as a reliable annotator, ensuring high-quality synthetic data in long-context scenarios.
[…]
En.QA Zh.QA
YaRN-Mistra 9.55 16.98
Kimi-Chat 16.52 18.62
Claude 2 11.97 10.53
GPT-4 22.22 23.06
LLaMA-3.1-8B-Instruct (Base) 27.11 29.77
LongPAI (ours) 32.46 32.38
Table 4: Ablation study on Supervised CoT Reasoning. Symbol § represents LongPAI without supervised CoT reasoning during fine-tuning. Average Scores (0-100) are evaluated by GPT-4-Turbo. Green highlights the remarkable improvements over the base GPT-4o-mini, while Red indicates a decline. Abbreviations: S.L. (Spotlight Locating), Comp. (Comparison), Clust. (Clustering), and Chain. (Chain of Reasoning).
Method S.L. Comp. Clust. Chain. Overall
$\mathtt{Set1}$ (10K-50K)
LLaMA-3.1-8B 89.13 72.33 31.77 74.0 60.50
LongPAI§ 93.39 68.67 79.38 83.80 79.82
LongPAI 85.00 90.00 69.12 98.30 81.58
$\mathtt{Set2}$ (50-100K)
LLaMA-3.1-8B 80.58 51.11 27.39 75.12 51.13
LongPAI§ 82.12 40.40 11.24 50.25 38.11
LongPAI 92.88 69.40 67.57 87.38 75.49
$\mathtt{Set3}$ (100K-200K)
LLaMA-3.1-8B 63.38 36.04 20.28 62.49 40.45
LongPAI§ 59.58 30.43 5.47 37.43 29.46
LongPAI 88.00 50.07 55.01 84.03 65.10
$\mathtt{Set4}$ (200K-250K)
LLaMA-3.1-8B 40.74 35.85 19.77 28.73 30.88
LongPAI§ 43.37 22.00 5.60 17.20 22.14
LongPAI 77.89 62.00 48.00 51.33 60.36
4.3 Discussion
Ablation Study on Supervised CoT Reasoning. To further analyze the impact of supervised CoT reasoning, we fine-tune the base LLaMA-3.1 model on LongFinanceQA while excluding CoT reasoning steps from the augmented answers, resulting in a new model, LongPAI§. Unlike the original LongPAI, LongPAI§ directly predicts the final answer, skipping intermediate reasoning steps. Table 4 shows that LongPAI significantly outperforms LongPAI§ over different input lengths. This result strongly supports the hypothesis mentioned in Section 1, which argues that directly guiding models to generate brief answers without intermediate reasoning steps for long-context modeling will lead to suboptimal training (finding 1). Moreover, Table 4 presents several interesting finding. First, while LongPAI§ performs comparably to LongPAI on short content (10K–50K tokens), its performance declines significantly on longer content, which means CoT reasoning is necessary for long-context modeling (finding 2). Furthermore, LongPAI§ performs well in single-source questions (Spotlight Locating) but struggles with multi-source questions (Comparison, Clustering, and Chain of Reasoning) as input length increases, demonstrating that CoT reasoning benefits complex long-context problem-solving (finding 3). In sum, all these findings reaffirm the importance of the reasoning capability for long-context modeling.
Comparison of Various Inference Frameworks. We compare PAI with two relevant inference frameworks: PAI- and RAG. PAI- is a variant of PAI that generates sub-questions directly, rather than first extracting properties and then generating sub-questions. Table 5 presents that PAI- generally outperforms the base GPT-4o-mini, except on $\mathtt{Set1}$. However, there is a gap between PAI- and PAI, highlighting the superiority of the Property Extraction Agent. Retrieval-Augmented Generation (RAG) Lewis et al. (2020) follows a two-step process: it first retrieves the top $K$ chunks relevant to a given query, and then uses retrieved chunks to generate an answer. Following Loong Wang et al. (2024), we adopt the BGE Chen et al. (2024a) as the embedding choice and set $K$ to 50, selecting from 5, 10, 30, and 50. Table 5 shows that RAG struggles with long-context problems, a conclusion also reached by Wang et al. (2024).
Table 5: Comparison of inference frameworks on the Loong Benchmark. Performance is evaluated by GPT-4-Turbo across four sets with different context sizes: $\mathtt{Set1}$ (10–50K), $\mathtt{Set2}$ (50–100K), $\mathtt{Set3}$ (100–200K), and $\mathtt{Set4}$ (200–250K). Green indicates an improvement over base GPT-4o-mini. Red denotes a decline.
Method $\mathtt{Set1}$ $\mathtt{Set2}$ $\mathtt{Set3}$ $\mathtt{Set4}$ Overall
GPT-4o-mini 65.05 50.63 45.41 31.61 49.25
          w/ PAI 78.69 69.62 68.95 58.12 69.58
          w/ PAI- 60.52 57.69 54.89 37.99 54.56
          w/ RAG 64.67 41.21 31.80 22.08 40.35
Efficiency Analysis on PAI. The PAI framework relies on multi-step inference, whereas LongPAI achieves results with a single inference step. We adopt the number of input tokens processed on the Loong benchmark (i.e., 1,600 samples) as a metric to compare the efficiency between PAI and LongPAI. In this comparison, PAI processes 3.53B tokens in total, determined by the GPT-4o tokenizer, whereas LongPAI requires only 112M tokens—less than 3% of PAI’s total. Despite PAI delivering the strongest overall performance, LongPAI stands out as the far more efficient approach, demonstrating a dramatic reduction in computational cost.
5 Conclusion
In this work, we introduce LongFinanceQA, a novel long-context synthetic dataset featuring reasoning-augmented answers for practical long-context questions. To generate these enriched answers, we develop the Property-driven Agentic Inference (PAI) framework, which simulates human-like reasoning through property extraction, retrieval, and summarization. Empirical analysis confirms the effectiveness of PAI and ensures the high quality of LongFinanceQA. Beyond PAI, we explicitly guide a language model in learning CoT reasoning using LongFinanceQA. A series of empirical results underscore the importance of reasoning-augmented long-context modeling.
Limitations
In this study, we investigated the effectiveness of supervised CoT reasoning using reasoning-augmented long-context synthetic data (LongFinanceQA). While effective, our approach has certain limitations. First, although the reasoning-enhanced LongPAI model demonstrates significant improvement in the financial domain, its ability to generalize to broader long-context scenarios remains uncertain. To address this, our future work will explore the impact of diverse data sources and different data scales. Second, while the PAI framework achieves strong performance on the Loong benchmark, it still relies on human-crafted prompts to guide agents in maintaining structured reasoning. Therefore, we plan to explore inference methods that enable autonomous reasoning with long-context inputs, minimizing reliance on extensive human-crafted prompts.
Appendix A Appendix
A.1 Trend of Long-Context LLMs
In this section, we present the context window sizes of closed-source and open-source LLMs in Figure 5. It indicates the gap between closed-source and open-source LLMs becomes smaller. In particular, the LLMs listed in Figure 5 include T5 Raffel et al. (2020), GPT-3 Brown et al. (2020), Codex Chen et al. (2021), T0 Sanh et al. (2022), Anthropic Priyanshu et al. (2024), InstructGPT Ouyang et al. (2022), CodeGen Nijkamp et al. (2023b), PaLM Chowdhery et al. (2023), U-PaLM Tay et al. (2023), Flan-T5 Chung et al. (2024), GPT-3.5 Turbo, LLaMA Touvron et al. (2023), GPT-4 Achiam et al. (2023), Claude 1.3, CodeGen2 Nijkamp et al. (2023a), PaLM2 Anil et al. (2023), Claude 2, LLaMA-2 Touvron et al. (2023), Qwen Bai et al. (2023), Kimi-Chat-200K, Yi-34B Young et al. (2024), Falcon 180B Almazrouei et al. (2023), Gemini Ultra Team et al. (2023), Mixtral Jiang et al. (2024), Qwen1.5 Qwen Team (2024), Gemini 1.5 Pro, Claude 3.0 Anthropic Team (2024), Kimi-Chat-2M, GPT-4 Turbo Achiam et al. (2023), LLaMA-3 Dubey et al. (2024), GPT-4o Hurst et al. (2024), Qwen2 Yang et al. (2024a), GPT-4o-mini Hurst et al. (2024), Mistral NeMo Mistral AI team (2024b), LLaMA-3.1 Dubey et al. (2024), Mistral Large 2 Mistral AI team (2024a), o1, Qwen-2.5 Yang et al. (2024b), Gemini 2.0 Pichai et al. (2024), DeepSeek-V3 Liu et al. (2024a), DeepSeek-R1 Guo et al. (2025), Qwen-2.5-1M Yang et al. (2024b), and o3-mini.
A.2 Prompts of Property-driven Agentic Inference (PAI)
In this section, we present the prompts for the three agents in the PAI framework: the property extraction agent, the property-based retrieval agent, and the summarization agent.
Property Extraction Agent. To extract properties, we employ the function-calling API of GPT-4o-mini to selectively retrieve the relevant metric and its corresponding subject from a given query. Figure 4 illustrates the property extraction process using function calling. In addition, we incorporate domain-specific examples within the function call to enhance accuracy across different domains. For instance, in the finance domain, we use “profit”, “revenue”, and “debt” as metric examples, while the “financial document title” serves as a subject example. In the legal domain, “verdict” represents the metric, and “legal judgment” serves as the subject. Similarly, in the academic domain, “reference” and “citation” are used as metric examples, with “paper title” as the subject.
Figure 4: Function calling details of property extraction agent (top) and property-based retrieval (bottom).
Figure 5: The trend of the context window size of recent large language models. The green dots refer to open-source LLMs, while the red dots denote closed-source LLMs.
Property-based Retrieval Agent. After obtaining the metric and its corresponding subject, we generate a sub-question in the format: “What was the <metric> of the <subject>?”. Then, the long-context input document is divided into a list of 1024-token chunks. Each chunk is evaluated to determine its relevance to the sub-questions, as shown in Figure 4 (bottom). After that, each sub-question is assigned relevant chunks. Next, we pack these relevant chunks and generate a sub-answer to the corresponding sub-question. Thus, this agent functions similarly to RAG Lewis et al. (2020).
Summarization Agent. Given the original query, the summarization agent summarizes a conclusion based on the sub-answers generated by the property-based retrieval agent.
Figure 6: The histogram of input token length in the LongFinanceQA (left). The proportion of single-source and multi-source QA tasks in the LongFinanceQA (right).
Figure 7: The pipeline of long-context synthetic question generation for the LongFinanaceQA dataset.
A.3 LongFinanceQA Dataset Statistics
Figure 6 presents a histogram of input token lengths in the LongFinanceQA dataset, which generally range from 50K to 250K tokens. Most input documents contain approximately 200K tokens. At the same time, we provide the proportion of single-source and multi-source QA pairs in the proposed LongFinanceQA dataset, where single-source QA pairs make up 45.6%, while multi-source QA pairs account for 54.4%.
Figure 8: Case study on two representative long-context QA problems.
A.4 Synthetic Question Generation
Figure 7 illustrates the procedure of synthetic Question generation. First, We create a financial metric pool with key metrics like profit, revenue, and cash flow (represented as the grey box). Then, we randomly select metrics and metadata of financial reports (e.g., company name and year) to generate long-context questions requiring single- or multi-source evidence (represented as the blue box). In addition, the types of multi-source questions are inspired by Loong Wang et al. (2024) and include comparison, clustering, trend analysis, etc.
A.5 More Empirical Results
Case Study on Practical Long-Context Problems. As shown in Figure 8, we present two representative long-context QA examples. Both questions are particularly challenging, as they require models to consider multiple pieces of evidence from the input text. Specifically, the case study includes the user query (white box), ground truth (yellow box), predictions from the base LLaMA-3.1-8B-Instruct model Dubey et al. (2024) (white box), and predictions from our LongPAI model (green box). In addition, incorrect predictions are highlighted in red, while correct ones are marked in green. The results demonstrate the superiority of the proposed LongPAI model, which benefits from supervised CoT reasoning. Furthermore, Figure 8 illustrates that LongPAI offers significantly greater interpretability than its base model.
A.6 Full Performances
Table 6 presents the full results of the Table 2 in the main manuscript. Also, we involve more popular LLMs Yang et al. (2024b); Guo et al. (2025) into the Table 6.
Model Context Spotlight Locating Comparison Clustering Chain of Reasoning Overall
Length AS PR AS PR AS PR AS PR AS PR
$\mathtt{All\ Set}$ (10K-250K)
DeepSeek-R1-Qwen-32B 128K 53.66 0.46 52.19 0.39 39.76 0.17 65.15 0.51 49.92 0.34
Qwen2-72B-Instruct 128K 59.80 0.47 61.12 0.43 34.32 0.06 74.68 0.50 53.20 0.32
Qwen2.5-72B-Instruct 128K 71.07 0.63 59.14 0.41 38.23 0.08 81.09 0.60 57.36 0.36
LLaMA-3-8B-Instruct-262K 262K 58.60 0.41 33.12 0.16 20.04 0.01 35.10 0.09 34.41 0.15
GLM4-9B-Chat 1000K 72.69 0.60 49.31 0.32 23.41 0.02 60.77 0.28 46.71 0.27
Qwen-2.5-14B-Instruct-1M 1000K 78.83 0.72 65.27 0.50 36.24 0.07 79.10 0.64 59.78 0.41
GPT-4o 128K 88.23 0.84 62.90 0.48 45.51 0.17 69.40 0.43 63.05 0.44
GPT-4o-mini (Base) 128K 70.90 0.59 59.37 0.38 36.33 0.06 79.58 0.58 56.50 0.34
GPT-4o-mini w/ PAI (ours) 128K 91.07 0.83 74.40 0.58 61.55 0.32 89.63 0.78 75.56 0.57
LLaMA-3.1-8B-Instruct (Base) 128K 67.84 0.56 47.12 0.30 24.62 0.02 63.63 0.34 45.88 0.26
LongPAI (ours) 262K 87.02 0.78 64.50 0.48 60.95 0.32 81.89 0.71 70.54 0.52
$\mathtt{Set1}$ (10K-50K)
DeepSeek-R1-Qwen-32B 128K 41.73 0.33 39.56 0.23 25.67 0.06 55.14 0.37 37.35 0.21
Qwen2-72B-Instruct 128K 88.04 0.83 89.33 0.83 43.00 0.17 93.50 0.80 71.46 0.57
Qwen2.5-72B-Instruct 128K 88.70 0.87 84.67 0.80 43.92 0.10 88.00 0.80 70.07 0.54
LLaMA-3-8B-Instruct-262K 262K 75.82 0.64 40.83 0.23 20.68 0.03 68.00 0.40 43.14 0.25
GLM4-9B-Chat 1000K 88.26 0.83 73.17 0.57 24.65 0.00 87.30 0.50 59.07 0.40
Qwen-2.5-14B-Instruct-1M 1000K 96.09 0.96 88.67 0.80 49.00 0.20 100.00 1.00 76.02 0.62
GPT-4o 128K 100.00 1.00 88.50 0.80 55.25 0.28 98.50 0.90 79.13 0.65
GPT-4o-mini (Base) 128K 97.39 0.96 81.83 0.67 46.50 0.15 100.00 1.00 73.35 0.56
GPT-4o-mini w/ PAI (ours) 128K 95.87 0.91 91.50 0.83 78.38 0.60 96.80 0.70 87.89 0.75
LLaMA-3.1-8B-Instruct (Base) 128K 89.13 0.87 72.33 0.60 31.77 0.05 74.00 0.60 60.50 0.45
LongPAI (ours) 262K 85.00 0.74 90.00 0.83 69.12 0.47 98.30 0.80 81.58 0.67
$\mathtt{Set2}$ (50K-100K)
DeepSeek-R1-Qwen-32B 128K 41.73 0.33 39.56 0.23 25.67 0.06 55.14 0.37 37.35 0.21
Qwen2-72B-Instruct 128K 74.88 0.65 68.60 0.51 40.70 0.09 87.00 0.72 62.38 0.41
Qwen2.5-72B-Instruct 128K 86.00 0.82 61.64 0.41 46.94 0.16 93.12 0.85 65.36 0.46
LLaMA-3-8B-Instruct-262K 262K 72.44 0.54 44.03 0.26 24.11 0.02 37.25 0.12 40.40 0.20
GLM4-9B-Chat 1000K 82.12 0.68 52.73 0.36 26.04 0.04 76.83 0.42 51.66 0.31
Qwen-2.5-14B-Instruct-1M 1000K 88.75 0.85 74.07 0.61 39.06 0.09 96.38 0.90 67.24 0.51
GPT-4o 128K 95.75 0.95 72.87 0.57 54.94 0.27 73.38 0.47 70.10 0.51
GPT-4o-mini (Base) 128K 82.65 0.70 58.77 0.41 40.89 0.08 92.50 0.78 61.61 0.40
GPT-4o-mini w/ PAI (ours) 128K 89.38 0.85 70.73 0.51 64.37 0.36 94.62 0.90 75.34 0.57
LLaMA-3.1-8B-Instruct (Base) 128K 80.58 0.70 51.11 0.33 27.39 0.02 75.12 0.47 51.13 0.30
LongPAI (ours) 262K 92.88 0.88 69.40 0.56 67.57 0.42 87.38 0.80 75.49 0.60
$\mathtt{Set3}$ (100K-200K)
DeepSeek-R1-Qwen-32B 128K 41.73 0.33 39.56 0.23 25.67 0.06 55.14 0.37 37.35 0.21
Qwen2-72B-Instruct 128K 47.00 0.33 48.07 0.27 25.79 0.00 69.37 0.34 42.98 0.20
Qwen2.5-72B-Instruct 128K 60.47 0.48 49.00 0.28 30.61 0.01 76.54 0.46 48.99 0.26
LLaMA-3-8B-Instruct-262K 262K 54.14 0.40 22.93 0.05 15.43 0.00 29.00 0.00 28.58 0.11
GLM4-9B-Chat 1000K 74.75 0.65 41.63 0.24 21.99 0.01 49.86 0.17 43.58 0.25
Qwen-2.5-14B-Instruct-1M 1000K 74.33 0.68 54.64 0.35 30.72 0.01 73.71 0.51 53.47 0.33
GPT-4o 128K 87.25 0.83 46.00 0.31 36.68 0.08 64.57 0.40 54.79 0.36
GPT-4o-mini (Base) 128K 63.05 0.53 53.48 0.24 29.80 0.01 72.37 0.46 50.03 0.26
GPT-4o-mini w/ PAI (ours) 128K 94.08 0.83 74.13 0.63 55.78 0.20 87.71 0.77 74.21 0.55
LLaMA-3.1-8B-Instruct (Base) 128K 63.38 0.47 36.04 0.19 20.28 0.00 62.49 0.26 40.45 0.20
LongPAI (ours) 262K 88.00 0.80 50.07 0.28 55.01 0.21 84.03 0.69 65.10 0.43
$\mathtt{Set4}$ (200K-250K)
DeepSeek-R1-Qwen-32B 128K 18.33 0.11 10.25 0.05 11.27 0.00 26.00 0.07 15.52 0.05
Qwen2-72B-Instruct 128K 41.85 0.19 39.75 0.15 29.17 0.03 41.67 0.07 37.23 0.11
Qwen2.5-72B-Instruct 128K 57.48 0.44 49.50 0.30 27.33 0.00 55.00 0.13 45.51 0.22
LLaMA-3-8B-Instruct-262K 262K 34.19 0.07 20.00 0.05 20.31 0.00 20.71 0.00 24.61 0.03
GLM4-9B-Chat 1000K 40.85 0.19 29.50 0.05 18.13 0.00 25.73 0.00 28.51 0.07
Qwen-2.5-14B-Instruct-1M 1000K 59.44 0.41 37.00 0.20 27.33 0.00 31.67 0.00 39.57 0.16
GPT-4o 128K 69.26 0.56 50.50 0.35 30.70 0.00 50.67 0.07 49.58 0.25
GPT-4o-mini (Base) 128K 48.37 0.26 50.00 0.30 28.70 0.00 48.33 0.07 42.30 0.15
GPT-4o-mini w/ PAI (ours) 128K 82.78 0.70 63.50 0.35 48.00 0.17 76.00 0.53 66.14 0.42
LLaMA-3.1-8B-Instruct (Base) 128K 40.74 0.30 35.85 0.20 19.77 0.00 28.73 0.00 30.88 0.13
LongPAI (ours) 262K 77.89 0.63 62.00 0.45 48.00 0.17 51.33 0.47 60.36 0.41
Table 6: Full results of Table 2. AS represents Avg Scores (0~100) and PR denotes Perfect Rate (0~1). Green indicates improvements compared to the base model; Red denotes a decrease compared to the base model.